# bibliography.bib

### v1

@@ 4 Vulnerable Activation Functions
4.1 ML Framework Functions
Thus far, our discussion of the functions used for activation functions have focused on low level mathematical used by ML framework developers inside of higher level framework code. A Tensorflow neuron with tanh() activation might not, for instance, directly return the output of a call to math library’s implementation of tanh(); there might be wrapping or processing of the result. This is of interest as we can also exploit input-dependent behavior at this higher framework or application level.
An example of this is the code to Tensorflow lite’s sigmoid() function (called Logistic() in the source), shown in LABEL:lst:sigmoid_in_tflite in Appendix B. We can see that on top of any side-channel leaks due to calls to std:exp(), a higher-level set of memory access patterns and instruction count cases are introduced by a conditional routing to different functions (or simply a return of 1) based on the input. In this way, this sigmoid() function can actually be attacked at two levels which we will describe in Section 6.2.
5 Methodology
First, we describe the overall concept for our attack (Section 5.1) against a neural network. We then explain in more detail how to attack the first layer (Section 5.2) and finally the subsequent layers (Section 5.3).
5.1 Attack Concept
The starting point for our attack is the above observation (cf. Section 4) that, unlike in cryptographic libraries, the math functions that underpin many ML activation functions have not been hardened to ensure they execute in constant time and without input-dependent branches and memory access patterns [^2] A classic example in the cache timing literature is Look-Up Tables that may be accessed in deterministic patterns based on the function inputs.. We believe this work has not been undertaken in the context of ML due to performance concerns or because security is seen as less of a focus.
$i_{1}$$w_{1}$$i_{2}$$w_{2}$$\Sigma$$f(\Sigma)$ Activation function $output$ Bias $b$
@@ 5.1 Attack Concept
Given that such code patterns take different numbers of instructions (and exhibit different memory access patterns) depending on the input, we can more generally assert that there is a correlation between different code paths and the execution trace. Though these differences might be subtle (e.g., a single instruction), they are measurable given a suitable side channel. Thus, if an attacker can determine which branch a function has taken, they learn something about the input or argument to that function.
As mentioned before, one function with these properties is expf(). Figure 3 shows the 11 different return cases of this function based on input and their associated CPU instruction counts (collected using gdb). We note that we found other optimised implementations of expf() that all use a similar algorithm and thus exhibit the same input-dependent memory access patterns. We focus here on the one used by SGX tlibc. For ease of discussion we labeled these cases using the names shown. For example, expf(4) puts us in case Normal (+)(+)( + ), whereas expf(110) results in the Overflow case. We define a threshold (shown in Figure 3 as dotted black lines) as an input value that is on the precise border of two output classes. For example, $88.72168$ is the value between Normal (+)(+)( + ) and Overflow.
We can also use knowledge of these regions and thresholds in the other direction: for example, if we know that an execution of expf() takes 17 CPU instructions (placing us somewhere in the Overflow region), we know the input to that function call must be greater than $88.72168$. We note that the cases for Overflow and Underflow are both early returns and as such their instruction counts are much lower than the other ‘normal’ returns which actually perform some sort of unique computation.
We can use these threshold values to leak information about the precise value of a partially controlled input. If we consider the scenario where an input to expf() is being multiplied by a hidden constant $c$ and we have access to the instruction count of the execution, we can reveal $c$ by finding (via e.g., a binary search) a threshold value between two instruction count classes and then dividing that value by our input. For example, if we find that the input $3.2164$ multiplied by $c$ leads to threshold value $88.72168$, $c$ must be $\nicefrac{{88.72168}}{{3.2164}}\approx 27.5842$. We therefore define $3.2164$ as a convergence point: an input to the system that causes a threshold value and thus a leak of the precise input to the function.
We now extend the above into a chosen input attack against a toy ML network. Consider an attack against a simple neural network with only a single neuron, as shown in Figure 2. The activation function is simply expf(). The goal is to recover $w_{1}$, $w_{2}$, and $b$. We are free to choose values for $i_{1}$ and $i_{2}$. We first note that the input to a neuron’s activation function is the sum of the products of various user-controlled inputs and their hidden weights, plus a hidden bias term ($\Sigma$ in Figure 2); this is similar to the case described previously except we have two hidden constants multiplied by two different inputs and an additional term, the bias, that is never multiplied with our inputs.
@@
[…]
@@
Finally, if we replace exp() above with a different function (such as sigmoid()) that also leaks threshold values, the only thing that changes is how to conduct the binary search: the recovery procedure is otherwise identical. Building on this concept, we will show how it is possible to recover the hidden parameters of more complex networks to a high degree of accuracy.
5.2 Full Attack on First Layer
@@
Continuing on from Section 5.1, we discuss how to attack the first layer of a production neural network. As before, the goal of this phase is to find a number of convergence point sets equal to the number of inputs plus one in order to have a system of equations we can solve to recover the weights and bias.
@@
If we can inject arbitrary values into the first layer, we iterate over every neuron in the first layer and recovery its hidden parameters in 2 phases: calibration and binary search.
5.2.1 The Calibration Phase
@@
In the first phase, we quickly produce sensible starting points and context for the second phase of the attack. We prepare a test input to the network that is all small non-zero values but set the first value of the array to a large value. If we run this test input and detect that the neuron overflowed, we know that this input is large enough to cause an overflow and that the weight between the first input and this neuron must be positive. If instead we produce an underflow, we know that the weight must be negative. If we are not in an overflow or underflow state (perhaps owing to a low magnitude weight). In this case, we simply multiply the large value in our test input by 10 and try again. We continue in this way until we find the sign and magnitude required to over/underflow this neuron from each input to the network. Note that we do not actually binary-search towards a threshold between two cases here; we merely attempt to reach an overflow/underflow case. In this case, we do not need to find a threshold for the instruction count leak to be informative.
@@
Recovering the signs for each weight is useful since some of the cases (most of the “normal” ones) are symmetrical about zero, meaning we cannot learn the sign of the input only based on the state. Recovering the sign before we begin the binary search means we always know exactly where we are in the state space. Recovering a value that overflows/underflows the neuron is useful to accelerate the attack. Though we could choose the maximum float value for the starting point of our binary search, the lower the initial guess, the fewer steps the search has to take. Searching "up" in the calibration step and then searching down in the second phase reduces the number of invocations needed.
5.2.2 The Binary Search Phase
@@
Armed with the information collected in the calibration phase, we continue the attack against the same neuron in phase 1. Now, we can binary search to find the desired number of convergence point sets for this neuron. For as many convergence sets as we need, we start by generating an input set that is made of $n$ locked random values (one for each input) except that the $i$’th input is instead set to the large positive value recovered during calibration. We feed the input into the network and note the return case by interpreting the CPU instruction counts for each neuron. Even if some of the weights have different signs, the overall $\Sigma$ should be dominated by the large input multiplied by its weight. If the neuron has overflowed, our input has caused the neuron to enter the state above the threshold between normal (+) and overflow and should search down. If the neuron has instead underflowed, we must be in the boundary between underflow and normal (-). If the input is negated before entering exp() as in sigmoid(), the direction to search should be reversed. We continue the binary search on that $i$’th input (searching back up if we fall below an underflow/overflow case in terms of magnitude) until we hit the desired search depth or our search does not change, because we reached the limit of decimal places for a floating point number. The tradeoff between depth and speed is shown in Table II.
@@
Using this algorithm, we can generate as many novel convergence sets as we require since the unchanging members of the input set are all randomised between inputs. It is worth noting that the system of equations does not produce usable results if the values on the right hand side are all the same threshold, so targeting multiple thresholds is required. One quick way to find more thresholds for the system of equations is to flip the signs of the inputs to ensure both the underflow-to-normal and overflow-to-normal thresholds are included in the solution set. Note however that it is not required to target the thresholds between underflow/normal and normal/overflow; these are just the easiest to find since we can force an under/overflow and then work our way up/down until we find the normal region. There are other thresholds within exp() (as shown in Figure 3) that can be exploited if they can be reached so long as we know the sign of the weight and thus in which precise state we are in.
@@
Once we have the requisite number of equations, we can solve the system of equations as described in Section 5.1 to recover the weights/biases for this first neuron. We then perform the exact same two-part algorithm on each remaining neuron until we solve the entire layer. The output from our recovery tool as it recovers one neuron is depicted in Appendix A.
5.3 Extending the Attack to Deeper Layers
@@
In order to extend the attack into subsequent layers, we must solve two key problems. First, how can we ‘unwrap’ the effect of already-solved early layers to insert arbitrary inputs beyond them, deeper into the network? And second, is it always possible to generate those values?
Unwrapping Solved Layers
@@
Our attack methodology for the first layer relies on being able to precisely craft a set of “target” in order to search towards threshold points. In order to be able to do this for deeper layers, we first have to “unwrap” all the recovered layers between the input layer and the layer we target. In the following set of equations, we show how to rearrange the normal operation of a neuron to isolate the input as opposed to the target.
$act(input*W+b)$ $=target$
$input*W+b$ $=act^{-1}(target)$
$input*W$ $=act^{-1}(target)-b$
[…]
$input$ $=(act^{-1}(target)-b)*W^{-1}$
Note that it is also possible to do this without matrix inversion for greater numerical efficiency, e.g., using numpy.linalg.lstsq(W, act_inv(target)-b).
@@
In these equations, $act$ and $actinv$ are known since we know the architecture of the network and $W$ and $b$ are known because we already solved this layer. Therefore, given a solved layer’s hidden parameters, a desired ‘target’ set of values, and the appropriate inverse activation function, we can calculate the input we must feed into the layer to produce a desired target as the output of that layer. Note also that the above assumes the activation function is invertible without a substantial loss of accuracy. Inverting these functions might lead to an inability to pass certain values forward, which we discuss below.
Expressive Power
@@
Even if we can algebraically calculate how to pass arbitrary values through a solved layer, subsequent layers are harder to attack because we are limited by the “expressive power” (see Table I) of the neurons before the target layer in terms of both magnitude and sign. There are two consequences to this.
@@
The first is some values may be impossible to generate in subsequent layers. For example, since the output of sigmoid() is always positive, baring negative biases, it would be impossible to send negative values into the next layer. Similarly, relu() can never produce a negative output. The second is that certain thresholds such as those between expf()’s overflow/underflow and normal cases have a high magnitude ($\approx 88.723{-}103.972$) relative to generally smaller magnitude weights and biases. sigmoid() has a very small output range ($0{-}1$) making it impossible to produce a large enough $\Sigma$ in the next layer to ‘reach’ one of these two thresholds unless either the network is very dense with small-magnitude positively weighted neurons or has some positive and large magnitude ($>1$) weights connected to it. This suggests denser networks are more expressive and therefore more vulnerable. In both of these cases, we rely on certain key hidden parameters, which means that certain networks might not be vulnerable to our attack. We however note that security guarantees should never assume a certain network architecture.
Attack Outline
@@
Nevertheless, extending the attack into deeper layers is possible with certain caveats. There is no need to focus only on the high magnitude thresholds we previously discussed between the underflow/overflow cases and the normal one; there are six in the range from $\approx$ -1.04–1.04 (see Figure 3). By searching anywhere within this region (e.g., near the subnormal case as in [Gongye20]) we can, as before, build convergence sets for our inputs and solve for the hidden parameters. Because we might pass through a low expressiveness activation function like sigmoid (output range from 0–1), which is likely being further reduced by multiplication with a weight with magnitude <1, we might not cross a single threshold point (the presence of the bias as a component of $\Sigma$ means that we might be shifted away from zero far enough that we cannot reach both sides of the subnormal case).
@@
6 Implementation and Orchestration
We trained several neural networks using “full” Tensorflow (in Python). In this section, we will first introduce a proof-of-concept attack on a case-study victim network with an expf() activation function. Next, we show how to attack a more realistic activation function and start to recover partial information from deeper layers.
6.1 Regression Model with expf()
Model
We first focus on a regression model to predict home insurance costs, which was trained on the “Medical Cost Personal Dataset” [choi18]. The network has three hidden layers. The first hidden layer contains 100 neurons that use the Exponential() activation function which is just a direct call to expf(). There are 11 input nodes, meaning that there are in total 1200 parameters (1100 weights and 100 biases) that we target. After the first layer, there is a second hidden layer of 10 neurons which uses a relu() based activation function. A final layer contains a single relu() neuron.
For this first network, we only focus on the first hidden layer. Despite that, we note that:

### v2

@@ IV-A Attack Concept
Given that such code patterns take different numbers of instructions (and exhibit different memory access patterns) depending on the input, more generally there is a correlation between different code paths and the execution trace. Though these differences might be subtle (e.g., a single instruction), they are measurable given a suitable side channel. Thus, if an attacker can determine which branch a function has taken, they learn something about the input or argument to that function.
As stated, one function with these properties is expf(). Figure˜2 shows the 11 different return cases of this function based on input and their associated CPU instruction counts (collected using gdb). We focus here on the expf() implementation from SGX’s tlibc, but we note that we found other optimised implementations all using a similar algorithm and thus exhibiting similar input-dependent memory access patterns. For ease of discussion we have labeled these 11 cases using the names shown in Figure˜2. For example, expf(4) results in case Normal $(+)$, whereas expf(110) results in Overflow. We define a threshold (shown in Figure˜2 as dotted black lines) as an input value that is on the precise border of two timing classes. For example, the value $88.723$ is between Normal $(+)$ and Overflow.
We can also use knowledge of these regions and thresholds in the other direction: for example, if we know that an execution of expf() takes 17 instructions (in the Overflow region), we know the input to that function call must be greater than $88.723$. We note that the cases for Overflow and Underflow are both early returns and as such their instruction counts are much lower than the other ‘normal’ returns which actually perform some sort of unique computation.
We can use these threshold values to leak information about the precise value of a partially controlled input. If we consider the scenario where an input to expf() is being multiplied by a hidden constant $c$ and we have access to the instruction count of the execution, we can reveal $c$ by finding (via e.g., a binary search) a threshold value between two instruction count classes and then dividing that value by our input. For example, if we find that the input $3.2164$ multiplied by $c$ leads to threshold value $88.723$, $c$ must be $\nicefrac{{88.723}}{{3.2164}}\approx 27.5846$. We therefore define $3.2164$ as a convergence point: an input to the system that causes a threshold value and thus a leak of the precise input to the function.
[…]
@@
We now extend the above into an attack against a toy ML network: consider a simple NN with only a single neuron, as shown in Figure˜3. The activation function is simply expf(). The goal is to recover $w_{1}$, $w_{2}$, and $b$. We are free to choose values for $i_{1}$ and $i_{2}$. We first note that the input to a neuron’s activation function is the sum of the products of various user-controlled inputs and their hidden weights, plus a hidden bias term ($\Sigma$ in Figure˜3); this is similar to the case described previously except we have two hidden constants multiplied by two different inputs and an additional term, the bias, that is never multiplied with our inputs.
@@
Here, we define $(i_{1},\,0)$, $(0,\,i_{2})$, and $(i_{3},\,i_{4})$ as convergence point sets, because each set of chosen inputs forces the activation function to a threshold value. Once we have enough equations (one for each unknown), we solve using linear algebra for the unknown weights and biases. Note that the thresholds do not have to be different, though the more thresholds are hit, the more robust the solution.
@@
Though it is possible to find convergence point sets like $(i_{1},\,0)$ and $(0,\,i_{2})$ above where all but one input is locked to zero and the remaining input is used to search for a threshold point, this is not required. In fact, by locking all but one input to some small magnitude random numbers (e.g., $i_{3}$ in the third equation) and searching on the single unlocked parameter ($i_{4}$), an arbitrary number of convergence point sets can be found. Finally, if we replace exp() above with a different function (such as sigmoid()) that also leaks threshold values, the only change is how to conduct the binary search; the recovery procedure is otherwise identical. Building on this concept, we will show how it is possible to recover the hidden parameters of more complex networks to a high degree of accuracy.
IV-B ML Framework Functions
Thus far we have largely assumed a low level mathematical definition for these functions. However, as discussed in Section˜III, a Tensorflow neuron with a tanh() activation might not directly return the output of a call to the underlying maths library’s implementation of tanh(); there might be wrapping or processing of the result at the framework level. This is of interest as we can also exploit input-dependent behaviour at this higher framework or application level. An example of this is Tensorflow lite’s sigmoid() function (called Logistic() in the source), shown in Appendix˜A. We can see that on top of any side-channel leaks due to calls to std:exp(), a higher-level set of memory access patterns and instruction count cases are introduced by a conditional routing to different functions (or simply a return of 0) based on the input. In this way, this sigmoid() function can actually be attacked at two levels which we will describe in Section˜V-B.
IV-C Capturing Traces with SGX-Step
To practically execute our attack, we used SGX-Step to collect the CPU instruction counts necessary to understand the state of our victim networks and apply the steps described above. This required that our networks run inside an SGX enclave. Given the constraints of SGX’s standard library replacement tlibc, we decided to employ Tensorflow Microlite, a software package for running (converted) Tensorflow lite models on microcontrollers, instead of “full” Tensorflow (which relies on a variety of system calls not easily provided by SGX). Our SGX enclave thus included a statically compiled Tensorflow Microlite library to perform inference. We converted our tensorflow models to Tensorflow lite and then further exported them as to a Tensorflow Microlite byte array which can be compiled directly into a C program. These byte arrays are added to the enclave code which allows us to perform “secure” inference inside SGX. We note that for practical deployments of SGX-protected ML, using such a reduced variant of Tensorflow would ease deployment, reducing code and thus Trusted Computing Base (TCB) size, while maintaining compatibility with Tensorflow-trained models. The underlying implementation [47] of exp() used by our network during inference comes from SGX’s libc alternative, tlibc.
We first profiled the execution of Tensorflow Microlite on an attacker-controlled debug enclave, so that we could understand execution flow by directly observing call traces via the interactive sgx-gdb debugger. An example can be seen in Appendix˜B. This provides an intuition for how Tensorflow Microlite performs inference and which functions and pages are accessed.
Figure 4: Inset of an SGX-Step instruction-granular page access trace showing the execution of the first layer of a NN with 16 neurons. Note that the intra-page instruction pointer (RIP) values (recovered by putting an enclave in debug mode) are displayed here only for reference purposes and are not used in our attack. We include them to show that knowing the current page provides enough information to interpret traces without precise RIP values.
We then developed an attack application, which starts and interacts with the victim enclave in production mode, i.e., without access to debug features. We use this to record page accesses which we later interpret into activation function states. Using SGX-Step, we log all page accesses within the production enclave (after a certain trigger page was accessed) and interpret the results as part of a post-processing step. Unlike sgx-gdb, SGX-Step only offers a relatively coarse-grained spatial resolution of 4 KiB memory pages, which prevents production enclave attackers from directly identifying the specific symbol (e.g., function or data) accessed by the enclave code. However, SGX-Step’s precise instruction-level temporal resolution enables the annotation of coarse-grained page-access traces with the exact number of instructions executed on each page. As illustrated in Figure˜4, this capability allows to accurately reconstruct the inference execution structure, even without knowledge of the lower 12-bit page offset of the secret enclave instruction pointer (ERIP). By comparing the dotted red line which depicts the true ERIP values available to sgx-gdb and the black line which shows the reduced page-level granularity, one can see how little is lost. It is worth nothing that each page can contain multiple functions and so it is difficult to assign a precise “meaning” to each page without context. The instructions executed on one page could pertain to two very different activities at different points in execution. However, the combination of coarse-grained page-access patterns along with the exact amount of instructions executed per page, annotated in black text, clearly suffices to identify individual function calls or data accesses. In practice, this means that even if multiple symbols are co-located on the same 4 KiB page, we can still interpret the logs. We also know precisely the length (or length range) of different runs at different points during the execution, i.e., we can also tell if a trace is invalid and immediately reject it.
By studying inference traces from the perspective of both sgx-gdb and SGX-Step (raw traces or visualizations like those shown in Figure˜4), clear patterns emerge. For example, the function ExpEval() is located on a different 4 KiB page than expf() (each 0x1000 bytes is a new page) and we can plainly interpret the movement between them as the higher level neuron code invoking the lower level expf() maths function 16 times (once for each neuron in the first layer of the model this plot was constructed from). The 18 instruction count long runs at the extreme left and right on page 0x2c000 seem to correspond to the start and end of a layer, which makes parsing each layer trivial. We note that the step counts provided by sgx-gdb and SGX-Step may vary slightly (but deterministically) due to the documented effects of macro-op fusion [38].
The final part of our tooling is a Python script and library which further abstracts and manages invocations of the enclave program. Because each neuron is accessed in order and each return occurs in sequence, it should be possible to monitor each neuron’s step count independently. In order to orchestrate the full attack, we use this script to interpret the log output by SGX-Step into a per-neuron state. Even though we receive the state for every neuron in each log, we only focus on one at a time. It takes about ten seconds to run each input through our SGX-Step setup and parse the trace. An example output of the running attack is shown in LABEL:lst:neuron_centric_attack in Appendix˜C.
IV-D Neuron-Centric Attack on First Layer
@@
Here we discuss how to convert the exploit discussed in Section˜IV-A into a full attack on the first layer of a production NN. As before, the goal of this phase is to find a number of convergence point sets equal to the number of inputs plus one in order to have a system of equations we can solve to recover the weights and bias.
@@
If we can inject arbitrary values into the first layer, we can conduct a chosen-input, neuron-centric attack pass by iterating over every neuron in the first layer and recovering their hidden parameters in 2 phases: calibration and binary search.
IV-D1 The Calibration Phase
Because the attack is a binary search to find hidden parameters of unknown value, can achieve dramatic speedups if we can find a reasonable upper bound to start the search from, rather than stepping down from the maximum possible float value. It is also useful to know the signs of the weights connected to our neurons. We can recover both pieces of information through the calibration step.
@@ IV-D1 The Calibration Phase
We prepare a test input to the network that is all small, non-zero values except for the first value which is a very large value. Next, we perform inference on this test input and consult the trace to see if the target neuron overflowed or underflowed. If it has, we know both that this input is already large enough to cause an overflow/underflow and the sign of the weight between the first input and this neuron. If the neuron has not overflowed or underflowed, (owing to a comparatively lower magnitude weight) we simply multiply the large value in our test input by 10 and try again. We continue in this way until we find the sign and magnitude required to over/underflow this neuron from each input to the network. Note that we do not actually binary-search towards a threshold between two cases here; we merely attempt to reach an overflow/underflow case. It is not necessary to find precise threshold points to learn the rough magnitude and sign of the inputs.
@@
Recovering the signs for each weight is useful since some of the cases (most of the “normal” ones) are symmetrical about zero, meaning we cannot learn the sign of the input only based on the state. Recovering the sign before we begin the binary search means we always know exactly where we are in the state space. Searching “up” in this calibration step gives us a sensible starting point to search down and thus reduces the number of invocations needed in the next phase.
IV-D2 The Binary Search Phase
@@
Armed with the information collected in the calibration phase, we continue the attack against the same neuron as in phase 1 by finding the desired number of convergence point sets. For as many convergence sets as we need, we start by generating an input set that is made of $n$ locked random values (one for each input) except that the $i$’th input is instead set to the large positive value recovered during calibration. We feed the input into the network and note the return case by interpreting the CPU instruction counts trace for the neuron. Even if some of the weights have different signs, the overall $\Sigma$ should be dominated by the large input multiplied by its weight. If the neuron has overflowed, our input has caused the neuron to enter the state above the threshold between normal (+) and overflow and should search down. If the neuron has instead underflowed, we must be in the boundary between underflow and normal (-). If the activation’s input is negated before being passed into exp() (as in sigmoid() ), the direction to search should be reversed. We continue the binary search on that $i$’th input, searching back up if we fall below an underflow/overflow case in terms of magnitude. Eventually the search ends either because the desired search depth was reached or because we reached the limit of precision for the underlying floating point maths.
@@
Using this algorithm, we can generate as many novel convergence sets as we require since the unchanging members of the input set are all randomised between inputs. It is worth noting that the system of equations does not produce usable results if the values on the right hand side are all the same threshold, so targeting multiple thresholds is required. One quick way to find more thresholds for the system of equations is to flip the signs of the inputs to ensure both the underflow-to-normal and overflow-to-normal thresholds are included in the solution set. Note however that it is not required to target the thresholds between underflow/normal and normal/overflow; these are just the easiest to find since we can force an under/overflow and then work our way up/down until we find the normal region. There are other thresholds within exp() (as shown in Figure˜2) that can be exploited if they can be reached so long as we know the sign of the weight and thus in which precise state we are in.
@@
Once we have the requisite number of equations, we can solve the system of equations as described in Section˜IV-A to recover the weights/bias for this first neuron. We then perform the exact same two-part algorithm on each remaining neuron until we solve the entire layer. The output from our recovery tool as it recovers one neuron is depicted in Appendix˜C.
IV-E Extending the Attack to Deeper Layers
@@
In order to extend the attack into subsequent layers, we must solve two key problems. First, how can we mitigate or ‘unwrap’ already-solved early layers to insert arbitrary inputs beyond them, deeper into the network? And second, is it always possible to generate those values?
@@
Unwrapping Solved Layers. Our attack methodology for the first layer relies on being able to precisely craft a set of “target” in order to search towards threshold points. In order to be able to do this for deeper layers, we first have to “unwrap” all the recovered layers between the input layer and the layer we target. In Appendix˜D, we show the simple linear algebra required to rearrange the normal operation of a neuron to isolate the input as opposed to the target.
@@
In these equations, actact and actinvactinv are known since we know the architecture of the network and $W$ and $b$ are known because they belong to a solved layer. Therefore, given these recovered parameters, a desired ‘target’ set of values, and the appropriate inverse activation function, we can calculate the input we must feed into the layer to produce a desired target as the output of that layer. Note this assumes the activation function is invertible without a substantial loss of accuracy; inverting these functions might lead to an inability to pass certain values forward, which we discuss below.
@@
Expressive Power. Even if we can algebraically calculate how to pass arbitrary values through a solved layer, subsequent layers are harder to attack because we are limited by the “expressive power” (see Table˜I) of the neurons before the target layer in terms of both magnitude and sign. There are two consequences to this.
@@
The first is some values may be impossible to generate in subsequent layers. For example, since the output of sigmoid() is always positive, baring negative biases, it would be impossible to send negative values into the next layer. Similarly, relu() can never produce a negative output. The second is that certain thresholds such as those between expf()’s overflow/underflow and normal cases have a high magnitude ($\approx 88.723{-}103.972$) relative to generally smaller magnitude weights and biases. sigmoid() has a very small output range (010{-}1) making it impossible to produce a large enough $\Sigma$ in the next layer to ‘reach’ one of these two thresholds unless either the network is very dense with small-magnitude positively weighted neurons or has some positive and large magnitude (>1>1) weights connected to it. This suggests denser networks are more expressive and therefore more vulnerable. In both of these cases, we rely on certain network properties. Though this means that certain networks might not be vulnerable to this method, security guarantees should never rest on architectural decisions.
@@
Attack Outline. Nevertheless, extending the attack into deeper layers is possible with certain caveats. There is no need to focus only on the high magnitude thresholds we previously discussed between the underflow/overflow cases and the normal one; there are six in the range from $\approx$ -1.04–1.04 (see Figure˜2). By searching anywhere within this region (e.g., near the subnormal case as in [27]) we can, as before, build convergence sets for our inputs and solve for the hidden parameters. Because we might pass through a low expressiveness activation function like sigmoid() (output range from 0–1), which is likely being further reduced by multiplication with a weight with magnitude <1, we might not cross a single threshold point (the presence of the bias as a component of $\Sigma$ means that we might be shifted away from zero far enough that we cannot reach both sides of the subnormal case).
@@
IV-F Estimating Attack Performance
We note that the number of executions to perform our attack is dependent on the architecture of the network (and therefore the number of parameters to recover) but also some factors the attacker can control such as the depth of the binary search. Naively, the total number $Q$ of queries to recover the first layer can be computed as: $Q=\,P\,\cdot\,N\,\cdot S$, where $P$ is the number of hidden parameters (in the first layer, e.g., one weight per input plus one for the bias) to solve for, $N$ is the number of neurons in the first layer, and $S$ is the maximum number of steps in the binary search, minimised by the calibration process described in Section˜IV-D1.
The $P$ value correlates to the minimum number of equations we require to solve a system of equations: one for each unknown. We note that we can compensate for noise or other errors by adding extra equations to produce better solutions at the cost of more queries. In testing, finding 3 additional convergence sets helped reduce the maximum error by 28% at depth 25 compared to an attack where precisely $P$ convergence sets were generated (this is visualised in Appendix˜K). It might be possible with future study to detect particularly weak systems of equations analytically and dynamically add extra convergence sets to improve the results.
V Experimental Evaluation
Here we introduce and share the results of our attack against three PoC networks trained on different datasets in “full” Tensorflow (in Python) and then converted to Tensorflow Microlite using the process described in Section˜IV-C.
V-A Minsurance: A Regression Model using expf()
Model. We first focus on a regression model to predict home insurance costs, which was trained on the “Medical Cost Personal Dataset” [53]. The network has three hidden layers. The first hidden layer contains 100 neurons that use the Exponential() activation function which is just a direct call to expf(). There are 11 input nodes, meaning that there are in total 1200 parameters (1100 weights and 100 biases) that we target. After the first layer, there is a second hidden layer of 10 neurons which uses a relu() based activation function. A final layer contains a single relu() neuron.
Recovery. For this first network, we only focus on the first hidden layer. Despite that, we note that: